OpenStudy (anonymous):

differentiate f(x)=cos(3x)sin^2(4x)

OpenStudy (anonymous):

is the answer -sin(3x)sin(8x)

zepdrix (zepdrix):

Hmmm did you forget to apply the product rule? <:o

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

Is the correct answer then (-sin(3x))(sin^2(4x))+(cos(3x))(sin(8x))

zepdrix (zepdrix):

\[\large f(x)=\cos(3x)\sin^2(4x)\]\[\large f'(x)=(\cos(3x))'\sin^2(4x)+\cos(3x)(\sin^2(4x))'\] \[\large = (-3\sin(3x))\sin^2(4x)+\cos(3x)(2\sin(4x)\cos(4x)(4))\] \[\large =-3\sin(3x)\sin^2(4x)+8\cos(3x)\sin(4x)\cos(4x)\] \[\large =-3\sin(3x)\sin^2(4x)+4\cos(3x)\sin(8x)\] Hmm I think you forgot the 4 that comes out from differentiating the innermost function of the squared term. Looks good besides that though! :)

OpenStudy (anonymous):

thanks

zepdrix (zepdrix):

Oh and you missed the 3 on the first term also, dont forget to do the chain rule multiple times if you need to :D Multiply by (3x)' on the first term, and (4x)' on the second Understand where those come from?

OpenStudy (anonymous):

yes i see it now thanks i knew that i was forgetting something